# ARKit LabelMaker: A New Scale for Indoor 3D Scene Understanding

### v1

@@ Abstract
The performance of neural networks scales with both their size and the amount of data they have been trained on. This is shown in both language and image generation. However, this requires scaling-friendly network architectures as well as large-scale datasets. Even though scaling-friendly architectures like transformers have emerged for 3D vision tasks, the GPT-moment of 3D vision remains distant due to the lack of training data. In this paper, we introduce ARKit LabelMaker, the first large-scale, real-world 3D dataset with dense semantic annotations. Specifically, we complement ARKitScenes [4] dataset with dense semantic annotations that are automatically generated at scale. To this end, we extend LabelMaker Weder et al. [33], a recent automatic annotation pipeline, to serve the needs of large-scale pre-training. This involves extending the pipeline with cutting-edge segmentation models as well as making it robust to the challenges of large-scale processing. Further, we push forward the state-of-the-art performance on ScanNet and ScanNet200 dataset with prevalent 3D semantic segmentation models, demonstrating the efficacy of our generated dataset.
@@ 1 Introduction
Recent progress in deep learning has been mostly focused on language [25, 26, 5] and 2D image generation [27]. Because for these two modalities, there is vast amount of training data available on the web. For text and image generation, you can simply scrape the internet for all available data and train your model in an auto-regressive (in the case of language) or diffusion process (image generation), where the supervision signal comes from self-supervision that does not require any labelling. This type of training revealed surprising properties [34] when scaling it to billions of data points in both text and image generation and leads to unseen performance gains enabling completely new use cases. Yet, this type of training is not available for scene understanding as objects and other primitives have to be classified in separate categories. This usually requires ground-truth annotations for the training data. While there are efforts to relax this requirement through self-supervision [12, 42] or open-set scene understanding [17, 36], state-of-the-art methods [36, 19] all have some form of direct supervision during training. Thus, annotated data is required for learning these tasks. Yet, creating datasets of similar scales as used in language and image generation is far from trivial. In this paper, we contribute the largest 3D real-world indoor semantic dataset and investigate the following questions: Is there a benefit from preferring real-world data over synthetic data? How can the labeling effort be reduced? Do current models profit from more real-world data?
To answer these questions, we make use of ARKitScenes [4], a large collection of RGB-D trajectories manually captured with consumer tablets. While these trajectories are annotated with sparse object bounding boxes, they are not complete enough to train competitive models for scene understanding and lack dense labels that could potentially serve as supervision for semantic segmentation models. Therefore, we supplement this dataset with dense semantic labels that we automatically generate using an automated pipeline. This allows us to create the first large-scale, dense 3D semantic segmentation training dataset that can be used to (pre-)train any 3D semantic segmentation model. To demonstrate the value of these vast yet imperfect annotations, we use these labels to re-train different models and extensively evaluate them on popular 3D semantic segmentation benchmarks.
@@
More specifically, we build on top of the recent LabelMaker [33] pipeline, which we extend into LabelMakerv2 with more and updated base models, a more general input data structure, scripts to deploy the pipeline onto large clusters through docker or SLURM. Using this pipeline, we process the entire ARKitScenes dataset, which takes 48000 GPU hours on Nvidia 3090 GPUs. Further, we scale this approach beyond ARKitScenes by integrating commonly used scanning software into the pipeline. This allows to robustly generate automatic labels for any scan acquired with ubiquitously available mobile devices. This opens up a road to potentially reach so far unseen scales in 3D datasets for scene understanding. The automatically generated ARKitScenes labels are used to pre-train the currently most-used 3D segmentation methods, MinkowskiNet [6] and PTv3 [36]. We find that without labeling any data manually, extending the scale of real-world training data improves the performance of both models on multiple benchmarks, or achieves the same performance as with even larger synthetic data.
@@
In summary, we answer the following research question: does large-scale pre-training with automatic labels show similar trends as it does for language and image generation tasks? We do that by making the following key contributions:
  • Improving LabelMaker [33] to LabelMakerV2 by incorporating new base models and improving robustness to large-scale scenes.
@@
  • Generating the largest existing real-world dataset with dense semantic annotations on 186 classes.
  • We push forward the state-of-the-art performance of point transformer on ScanNet and ScanNet200 by a large margin which our gerenated dataset.
The ArkitLabelMaker dataset and LabelMakerV2 pipeline code is available at labelmaker.org.
@@ 2 Related Works
Datasets for 3D semantic segmentation. 3D semantic segmentation is to classify each point in a 3D point cloud into a set of predefined semantic categories. Prominent datasets for training and evaluation include ScanNet [10]/ScanNet200 [29], and the Stanford 3D Indoor Scene Dataset [1] (S3DIS), which comprises 6 large-scale indoor areas with 271 rooms. Both datasets includes RGB-D frames captured from real-world. In addition to real-world datasets, Structured3D [41] offers a photo-realistic synthetic dataset with 3.5K house designs, and Replica [31] provides 18 high-quality reconstructed scenes. ARKitScenes [4] is the most extensive collection of indoor scenes to date, featuring 5047 captures of 1661 unique scenes. RGB-D data is captured with Apple LiDAR scanner. High-quality surface reconstruction and the bounding box for object detection are also provided. However, dense annotations for semantic or instance segmentation is absent from this dataset. Therefore, it cannot be directly used in training models for 3D semantic segmentation.
LabelMaker Weder et al. [33] is an automatic semantic segmentation annotation pipeline that consolidates outputs from state-of-the-art 2D segmentation models with an additional feature for translating frame-wise 2D labels into cohesive 3D point cloud labels. In this work, we employ an enhanced version of LabelMaker to generate dense semantic segmentation annotations for ARKitScenes.
3D semantic segmentation models. The neural network architecture for processing 3D input data can be classified into three main categories: voxel-based, point-based, and transformer-based methods. Voxel-based methods transform points into fixed-sized voxel grids before passing them through the neural network. MinkowskiNet [6] is one of the most well-known model. Mix3D [21] enhances MinkowskiNet through effective 3D data augmentation techniques. PonderV2 [42] explores self-supervised learning from RGB-D data to improve the performance of the MinkowskiNet architecture. Point-based methods includes [23, 24, 16, 15, 38, 32, 11, 3]. However, there is a recent shift from models based on point-wise convolutions to point-based transformer models [22, 14, 40, 30]. Notable examples include Point Transformer [40], PTv2 [35], and PTv3 [36], which are developed towards better efficiency and scaling ability for 3D inputs. Point Prompt Training [37](PPT) introduces a novel training paradigm enabling the simultaneous training of multiple datasets with diverse label spaces. Combining PTv3 with PPT achieves state-of-the-art performance on the ScanNet/ScanNet200 semantic segmentation benchmark.
In this paper, we address the main limitation of existing datasets for 3D semantic segmentation: their limited size. We suspect that this limited size negatively impacts the performance of commonly used models as their performance is limited without additional training data available.
@@
3.1 LabelMaker
As we build on top of LabelMaker [33], we briefly review the essential steps of the pipeline proposed by Weder et al.. LabelMaker is an automatic labelling pipeline for 2D and 3D semantic annotation. As shown in [33], it generates labels that are on par with the human annotators in [9]. It automatically generates these labels by exploiting an ensemble of base models that predict semantic maps for every input frame in an RGB-D trajectory. Since the base models predict segmentation in different label spaces based on their training data, they are then mapped to a unified label space. Only through this mapping, the different base models can be used in a subsequent ensemble. Thus, [33] defined a mapping from every label space into a carefully curated label space based on wordnet synkeys. After mapping all base model predictions into the unified label space, the individual predictions are aggregated into one single consensus for every frame of the RGB-D trajectory. As the RGB-D trajectory provides multi-view information of the scene, the individual frames can be further denoised by lifting them into 3D and aggregate the point-wise predictions across the different frames. The final labels can either be directly used as 3D labels for 3D semantic segmentation or projected into 2D and be used for training or evaluating 2D semantic segmentation models. In this paper, we improve this pipeline to robustly scale to large-scale datasets and show its benefit to pretraining 3D semantic segmentation models. In the following, we describe the improvements and the scaling in more detail.
@@ 3.1 LabelMaker
Figure 1: Dependency graph of the LabelMakerv2 pipeline. Our LabelMakerv2 pipeline has a clear dependency structure that has to be handled in the distributed processing of the data. This has to be especially respected when recovering from job failure. There, our recovery strategy checks for unfinished jobs in the dependency graph before submitting any new jobs to avoid unnecessarily wasting compute resources. The boxes with thick yellow frame donotes visualizable tasks. These are used during inspection and job quality assurance.
3.2 Improving Labelmaker to make it scale to ARKitScenes [4]
@@
While LabelMaker [33] presented an automatic labelling tool that produces annotations on par with human annotators, we enhance the pipeline with two changes to further improve its performance that is necessary to robustly generate high-quality annotations for large-scale datasets. The complete pipeline is shown in Figure 1.
@@
Integrating Grounded-SAM LabelMaker [33] used several state-of-the-art base models in its esemble. Yet, they did not leverage the potential of Segment Anything (SAM) [13], a 2D segmentation model that was trained on a large-scale datasets and robustly generalizes to many scenarios. As we want to scale LabelMaker to any environment, our aim is to integrate this prior into the pipeline. However, it is in not straight forward to efficiently use this model for semantic segmentation. To this end, Grounded SAM combines Grounding DINO [18] with SAM [13]. Grounding DINO locates instances’ bounding boxes given semantic labels or natural languages, while the later model generates high quality segmentation masks for these bounding boxes. We integrate this model by adapting it to LabelMaker’s unified labelspace such that it can act as an additional vote in the ensemble.
@@
Aligning to gravity For ideal performance, many semantic segmentation models require alignment of the gravity direction with the coordinate system of the data they were trained on. Yet, large-scale data is not aligned with gravity by default. For example in ARKitScenes, occasional phone rotation leads to inconsistencies in the orientation of 2D images. Passing those rotated images to LabelMaker’s base models results in decreased performance and misclassifications (e.g. mistaking the floor with the ceiling). Therefore, we project sky direction, which corresponds to the z-axis of the pose coordinate system from ARKit (that uses the IMU), onto each 2D frame. Then, we compute the angle between sky direction and upward direction $\alpha$. Given this angle, we rotate the image by $k\cdot\frac{\pi}{2}$, where $k=\arg\min_{s}(|s\frac{\pi}{2}-\alpha|)$ to make the sky direction roughly upward, and rotate the predicted segmentation back to its original orientation after passing it through model to align it with its coordinate system.
Optimizing compute resource scheduling. We optimize the code to deploy each individual piece of the pipeline of Figure 1 as individual jobs to a GPU cluster, with SLURM as a dependency manager between the pipeline pieces. To optimize the overall execution time, it is therefore important to be able to estimate the processing time of each piece of the pipeline at the point of job submission. ARKitScenes contains scenes of a wide range of sizes, spanning from a minimum of 65 frames to a maximum of 13796, and different parts of the pipeline scale differently with increasing scene size. To figure out the minimum resources requirements, we select 11 scenes of varied sizes uniformly distributed within the minimum and maximum range and record their resources usage. While most jobs are not sensitive to scene size and can suffice with a fixed resource allocation, the base models exhibit greater sensitivity to scene size. We interpolate resource needs with respect to scene size and summarize the empirical numbers in the Appendix. Through this, we ensure that we request minimal-required resources, so that we have lowest job waiting time and less idle compute power.
3.3 Scaling beyond existing datasets
In this paper, we demonstrate the effectiveness of the automated labels generated for a large-scale dataset for pre-training 3D semantic segmentation models. Yet, ARKitScenes is still limited in terms of variance in scene type. Through modern mobile devices, RGB-D scanning is ubiquitously available and a excellent source for 3D data. Yet, it has not been possible to exploit this data for 3D semantic segmentation without expensive human annotation. While LabelMaker [33] provided an automatic annotation pipeline, it was not readily available for data provided by mobile devices. To overcome this limitation, we integrated a well-established scaninng platform for iOS (Scanner 3D) into our imporved LabelMaker pipeline. This allows anyone with an iOS device to collect data and automatically generate human-quality annotations that can either be used for training or evaluation of 3D semantic models at an unseen scale.
To this end, we built a integration of Scanner 3D into LabelMakerv2. This involves two essential steps. First, we align the coordinate system of Scanner3D with the coordinate system of LabelMakerv2 by transforming all individual frames into the LabelMakerv2 system. Second, we have to solve the challenge that does not provide direct access to the individual depth frames but only provides a dense reconstructed mesh. To overcome this hurdle, we render a depth map from the mesh for every camera view in the scanned trajectory. Therefore, we also align the mesh with the LabelMakerv2 coordinate system and use Open3D to render the corresponding depth maps. Together with the corresponding RGB images and camera poses, these depth maps are subsequently transformed into the LabelMakerv2 format such that the trajectory can be processed by LabelMakerv2.
@@ 4.1 Baselines
We evaluate the effectiveness of our ARKitScenes LabelMaker dataset on two distinct network architectures: MinkowskiNet [6] and Point Transformer [8, 35, 37, 36]. MinkowskiNet is the most established architecture. Many modifications [19, 42] have been proposed to MinkowskiNet and it is still the underlying architecture of most top-performing models in 3D Semantic Segmentation benchmarks. Point Transformer [36] is a very recently proposed architecture and the current state-of-the-art on the ScanNet and ScanNet200 benchmarks. Since transformers are in general known to profit from large-scale training data, we also train on this architecture. From these two architectures, we derive our three relevant baselines:
@@
Vanilla MinkowskiNet. This is the standard MinkowskiNet model based on [6], which most 3D semantic segmentation methods compare to. In this paper, we use the commonly used ‘Res16UNet34C‘variant of MinkowskiNet to guarantee fair comparison to all other baselines.
PonderV2 [42]. PonderV2 [42] is an unsupervised pre-training strategy for semantic segmentation. For large-scale training data, manual labelling of training data is infeasible and leaves essentially two alternatives: Automatic labelling of the training data, which is what we investigate in this work, or an unsupervised strategy that learns useful features on the data from a different loss. PonderV2 is the state-of-the-art method for unsupervised training. While the best performing ponder-based model combines different training datasets, model architecture modifications, and the unsupervised pretraining, our evaluation focuses on comparing the two pretraining paradigms.
@@
PointTransformerV3 (PTv3) [36]. Point Transformer V3 (PTv3) [36] is a recently proposed method that aims to speed up transformer architecture and scale up. Point Prompt Training (PPT) is a large-scale training paradigm aiming to jointly train multiple datasets of various label spaces together. This is a different approach than the LabelMaker [33] approach of translating label spaces and then training on a common label space. The combination of PTv3 and PPT achieves state-of-the-art results on ScanNet/ScanNet200 semantic segmentation benchmark.
4.2 Datasets and Metrics for Evaluation
ScanNet [9]. ScanNet comprises 1513 densely annotated scans across 707 distinct indoor scenes, totaling 2.5 million RGB-D frames. It stands as one of the most widely used and influential benchmark datasets for indoor 3D scene understanding. ScanNet is annotated by humans using the NYU40 label space and evaluated on a subset of 20 classes from NYU40.
ScanNet200 [28]. While only 20 classes are used in the ScanNet benchmark, the original dataset is annotated with many more classes. ScanNet200 [28] leverages these annotations and organizes them into a new benchmark with 200 classes that are of higher-resolution than the original Scannet classes. Given the large-number of different categories generated by our LabelMakerv2 pipeline, we also pre-train the models for this task and evaluate them on the ScanNet200 benchmark.
ScanNet++ [39]. ScanNet++ is a dataset of 460 high-resolution 3D indoor scenes with dense semantic and instance annotations, captured using a high-precision laser scanner and registered images from a DSLR camera and RGB-D streams. It focuses on long-tail and multi-labeled annotations. In its semantic segmentation benchmark, models are evaluated over 100 labels.
S3DIS [2]. S3DIS is a 3D semantic dataset containing 6 large-scale indoor areas from 3 different buildings, labeled with 13 semantic classes. We follow the practice of [36] and create a dataset with an effective size of 406. We only use this dataset during the training of the PTv3 [36] baseline.
Structured3D [41] is a large-scale indoor synthetic RGB-D dataset featuring 6519 training scenes and 1697 test scenes. It is annotated with a label space of 25 classes. Structured3D and S3DIS only used in PTv3+PPT joint training and we adopt pre-processed version of these two datasets from [36]. The actual dataset size used in training is different from the original, and we refer to the actual number in Table 1.
ARKit LabelMaker (Ours). This is the dataset generated with the our method described above. The resulting dataset contains 5019 trajectories, from which we take 4471 for training and 548 for validation according to the official train-val split provided by the original ARKitScenes [4] dataset. For every scene, we created 3D pointcloud associated dense semantic labels in the original LabelMaker wordnet label space (186 classes). We denote this dataset as ALC. Moreover, we utilized the provided mapping from wordnet label space to the ScanNet200 label space and mapped all labels to this space. This resulted in the our ALS200 dataset. We use this data to train the different baseline models described above. To increase efficiency and make the experimental settings comparable to previous studies, we perform down-sampling on 3D meshes to a voxel size of 2cm. Normal information is preserved and down-sampled simultaneously. Table 1 illustrates the scale of each dataset. ARKitScenes LabelMaker dataset is the largest diversely annotated real-world indoor semantic dataset.
Metrics. We follow the standard metrics of the ScanNet semantic segmentation task and compute the mean and per-class intersection-over-union ratio (IoU) on ScanNet/ScanNet 200.
Dataset #train #val #test real #label

### v2

@@ Abstract
Neural network performance scales with both model size and data volume, as shown in both language and image processing. This requires scaling-friendly architectures and large datasets. While transformers have been adapted for 3D vision, a ‘GPT-moment’ remains elusive due to limited training data. We introduce ARKit LabelMaker, a large-scale real-world 3D dataset with dense semantic annotation that is more than three times larger than prior largest dataset. Specifically, we extend ARKitScenes [4] with automatically generated dense 3D labels using an extended LabelMaker pipeline [37], tailored for large-scale pre-training. Training on our dataset improves accuracy across architectures, achieving state-of-the-art 3D semantic segmentation scores on ScanNet and ScanNet200, with notable gains on tail classes. Our code is available at labelmaker.org and our dataset at huggingface.
Guangda Ji1      Silvan Weder1      Francis Engelmann1,2      Marc Pollefeys1      Hermann Blum1,3
1ETH Zürich       2Stanford University       3University of Bonn & Lamarr Institute
@@ 1 Introduction
Recent advancements in deep learning on language [26, 27, 5] and 2D vision  [28, 29, 1] have made tremendous progress, primarily driven by the abundance of training data available on the web for these modalities. Scaling this type of large-scale training to billions of data points has revealed surprising properties [38] and resulted in unprecedented performance gains, enabling entirely new applications. However, this approach is not directly applicable to 3D scene understanding, where real-world 3D data lacks web-scale abundance and demands labor-intensive ground truth annotations.
While recent efforts aim to reduce this dependency through self-supervision [13, 49], distillation [23, 45], denoising [35], or open-set scene understanding [18, 40, 44, 33, 10], state-of-the-art 3D segmentation methods [40, 21, 31, 46] still rely on some level of direct supervision. Consequently, annotated data remains essential for learning these tasks, and constructing datasets of comparable scale to those in language and image generation remains a significant challenge. In this paper, we contribute the largest 3D real-world indoor semantic dataset and investigate key questions in 3D scene understanding: Is real-world data preferable to synthetic data? How can labeling efforts be minimized? Do current models benefit from increased real-world data?
num scenes01000200030004000050006000
  train
  val
  test
Datasets SizeARKit LabelMaker(186 classes)ScanNet/ScanNet200 [9, 30](20/200 classes)S3DIS [2](13 classes)ScanNet++ [43](100 classes)
Figure 1: Our LabelMaker annotation data creates the world’s largest real-world 3D scene annotation dataset.
To address these questions, we leverage ARKit-Scenes [4], a large-scale 3D indoor dataset consisting of 3D reconstructions and RGB-D frames captured with consumer tablets. Although these scenes are annotated with 3D object bounding boxes, they lack the per-point annotations necessary for training competitive 3D segmentation models. To overcome this limitation, we augment the dataset with per-point semantic labels created through an automated pipeline. This approach enables us to produce the significantly larger dataset compared to prior 3D semantic segmentation dataset (see Fig. 1). Our dataset is suitable for (pre-)training any 3D semantic segmentation model. To validate the effectiveness of these extensive yet imperfect annotations, we use them to re-train various models and conduct comprehensive evaluations on widely-used 3D semantic segmentation benchmarks [30, 9].
@@
More specifically, we build on top of the recent LabelMaker [37] pipeline, which we extend into LabelMakerV2 with more and updated base models, a more general input data structure, as well as deployment scripts for large clusters through docker or SLURM. Using this pipeline, we process the entire ARKitScenes dataset, which takes 48’000 GPU hours on Nvidia 3090 GPUs. We further scale our pipeline beyond ARKitScenes to arbitrary scenes by integrating the iOS app Scanner 3D into LabelMakerV2, enabling automatic annotation of scenes recorded with consumer iPhones. In experiments, we use our automatically generated ARKitScenes labels to pre-train the currently most-used 3D segmentation methods, MinkowskiNet [7] and PTv3 [40]. We find that without labeling any data manually, extending the scale of real-world training data improves the performance of both models on multiple benchmarks, or achieves the same performance as with even more synthetic training data.
@@
In summary, we answer the following research question: “Does large-scale pre-training with automatic labels show similar trends in 3D as it does for language and image tasks?" through the following key contributions:
@@
  • Generating the largest existing real-world 3D dataset with dense semantic annotations on 186 classes.
  • Improving over state-of-the-art PointTransformer on ScanNet200 by 2.1%, on tail classes by 5.5% mIoU.
  • Trade-off analysis between training in unsupervised settings, on synthetic data, and on auto-labeled real-world data providing guidance for future data scaling efforts.
@@ 2 Related Works
Datasets for 3D semantic segmentation.
3D semantic segmentation classifies each point in a 3D point cloud into a set of predefined semantic categories. Prominent datasets for training and evaluation include ScanNet [9]/ScanNet200 [30] consisting of 1.5k scenes, and the Stanford 3D Indoor Scene Dataset [2] (S3DIS), which comprises 6 large-scale indoor areas with 271 rooms. Both datasets include RGB-D frames captured in the real world. In addition to real-world datasets, Structured3D [48] is a photo-realistic synthetic dataset with 3.5K house designs, and Replica [32] provides 18 high-quality reconstructed scenes. ARKitScenes [4] is the most extensive collection of indoor scenes to date, with 5047 scans of 1661 unique scenes. RGB-D data is recorded with an Apple iPad equipped with a built-in LiDAR scanner. High-quality surface reconstruction and the bounding box for object detection are also provided. However, per-point annotations for semantic or instance segmentation are not included, so it cannot be directly used for training 3D segmentation models.
LabelMaker.
Weder et al. [37] is an automatic 3D semantic segmentation annotation pipeline that consolidates outputs from state-of-the-art 2D and 3D segmentation models with an additional feature for translating frame-wise 2D labels into consistent 3D point cloud labels. In this work, we employ an enhanced version of LabelMaker to create 3D semantic segmentation annotations for ARKitScenes.
3D semantic segmentation models.
Deep learning models for processing 3D input data can be classified into three main categories: voxel-based, point-based, and transformer-based methods. Voxel-based methods transform points into fixed-sized voxel grids before processing them, such as the popular MinkowskiNet [7]. Mix3D [21] enhances MinkowskiNet through effective 3D data augmentation techniques. PonderV2 [49] explores self-supervised learning from RGB-D data to improve the performance of the MinkowskiNet architecture. Point-based methods includes [24, 25, 17, 16, 42, 34, 11, 3]. However, there is a recent shift from models based on point-wise convolutions to point-based transformer models [22, 15, 47, 31]. Notable examples include PointTransformer [47] and its successors PTv2 [39], and PTv3 [40], which are developed towards better efficiency and scalability. Point Prompt Training [41](PPT) introduces a novel training paradigm enabling the simultaneous training of multiple datasets with diverse label spaces. Combining PTv3 with PPT achieves state-of-the-art performance on the ScanNet/ScanNet200 semantic segmentation benchmark.
In this paper, we address a key limitation of existing datasets for 3D semantic segmentation: their small scale. We hypothesize that this constraint hampers the performance of commonly used models.
@@ 3 Method
3.1 LabelMaker Revisted
@@
Figure 2: Dependency graph of the LabelMakerV2 pipeline. Our LabelMakerV2 pipeline has a clear dependency structure that has to be handled in the distributed processing of the data. This has to be especially respected when recovering from job failure. There, our recovery strategy checks for unfinished jobs in the dependency graph before submitting any new jobs to avoid unnecessarily wasting compute resources. The boxes with thick green frame donotes visualizable tasks. These are used during inspection and job quality assurance.
As we build on LabelMaker [37], we briefly review its key steps. LabelMaker is an automatic pipeline for 2D and 3D semantic annotation, producing labels comparable in quality to human annotations [9]. It automatically generates semantic labels by leveraging an ensemble of base models to predict pixel-level semantics for each frame in an RGB-D trajectory. Since the base models predict segmentations in different label spaces (based on their training data), their predicted semantic labels are then mapped to a unified label space. Only through this mapping, the different base models can be used in a subsequent ensemble. Thus, [37] defined a mapping from every label space into a carefully curated label space based on wordnet synkeys [20]. After mapping all base model predictions to the unified label space, they are aggregated into a single consensus per frame of the RGB-D trajectory. This is the first stage of per-frame denoising. As the RGB-D trajectory provides multi-view information of the scene, the individual frames can be further denoised by lifting 2D predictions onto 3D points and performing per-point voting. The final labels can either be directly used as 3D labels for 3D semantic segmentation or projected into 2D and be used for training or evaluating 2D semantic segmentation models. In this paper, we improve this pipeline to robustly scale to large-scale datasets and show its benefit for pretraining 3D semantic segmentation models. In the following, we describe the improvements in more detail.
3.2 Improving LabelMaker for Scaling
@@
While LabelMaker [37] introduced an automatic labeling tool that produces annotations comparable to human annotators, we enhance the pipeline with two modifications to further improve its performance, ensuring robust high-quality annotation for large-scale datasets. The complete pipeline is shown in Figure 2.
Integrating Grounded-SAM.
@@
LabelMaker [37] employs several state-of-the-art base models in its ensemble, but does not utilize Segment Anything (SAM) [14], a 2D segmentation model trained on large-scale datasets that generalizes robustly across diverse scenarios. To scale LabelMaker to any environment, we aim to integrate this prior into the pipeline. However, efficiently leveraging this model for semantic segmentation is not straightforward. To address this, Grounded SAM combines Grounding DINO [19] with SAM [14]. Grounding DINO predicts instance bounding boxes based on semantic labels or natural language, while SAM generates high-quality segmentation masks for these boxes. We integrate this model by adapting it to LabelMaker’s unified label space, allowing it to contribute as an additional vote in the ensemble.
Without gravity alignment       With gravity alignment
Figure 3: Qualitative Evaluation of Gravity Alignment. LabelMaker annotation with and without gravity alignment. Without gravity alignment, floors may be misclassified as walls, walls as ceilings, as well as other orientation-dependent objects.
Aligning to Gravity.
@@
For optimal performance, many semantic segmentation models require the gravity direction to be aligned with the coordinate system used during training. However, large-scale datasets are not inherently gravity-aligned. For instance, in ARKitScenes, occasional phone rotations introduce inconsistencies in the orientation of 2D images. Passing these misoriented images to LabelMaker’s base models degrades performance and leads to misclassifications, such as confusing the floor with the ceiling (see Fig. 3). Therefore, we project the sky direction, corresponding to the z-axis of ARKit’s pose coordinate system (derived from the IMU), onto each 2D frame. We then compute the angle $\alpha$ between the sky direction and the upward direction. Given this angle, we rotate the image by $k\cdot\frac{\pi}{2}$, where $k=\arg\min_{s}(|s\frac{\pi}{2}-\alpha|)$ to align the sky direction roughly upward, and rotate the predicted segmentation back to its original orientation after inference to align it with its coordinate system.
Omission of NeuS.
In LabelMaker [37], an implicit surface model, NeuS [36] with a semantic head, is trained per scene as an optional 3D lifting and 2D denoising step. We omit this step due to its high computational cost. Moreover, NeuS optimizes its own scene geometry, which is poorly constrained in ‘inside-out’ scans of ARKitScenes, making point cloud label generation more complex than a simple coordinate lookup. Instead, we retain per-point voting from [37], as it proved the most stable in our initial exploration.
@@ 4.1 Baselines
We evaluate the effectiveness of our ARKitScenes LabelMaker dataset using two well-established and distinct network architectures: MinkowskiNet [7] and PointTransformer [8, 39, 41, 40]. MinkowskiNet remains the foundation of many top-performing models in 3D semantic segmentation benchmarks, with several modifications [21, 49] proposed to enhance its performance. PointTransformer [40], a more recent architecture, achieves state-of-the-art results on the ScanNet and ScanNet200 benchmarks. Given that transformers generally benefit from large-scale training data, we also train on this architecture. From these two architectures, we derive three relevant baselines:
Vanilla MinkowskiNet.
@@
This is the standard MinkowskiNet model based on [7], which most 3D semantic segmentation methods compare to. In this paper, we use the commonly used ‘Res16UNet34C‘variant of MinkowskiNet to guarantee fair comparison to all other baselines.
Mix3D [21]
is a data augmentation method for large-scale 3D scene segmentation that creates new training samples by merging two augmented scenes, effectively placing object instances into novel, out-of-context environments. This approach encourages models to infer semantics from local structures rather than relying on overall scene context. MinkowskiNet shows significant performance improvements when trained with Mix3D.
PonderV2 [49].
Addressing the scarcity of 3D annotations involves two strategies, unsupervised feature learning and automated pseudo-labeling. PonderV2 [49] represents the former, using neural rendering objectives to learn 3D features without semantic annotations. Our work explores the latter, generating large-scale pseudo-labels through automatic annotation. While PonderV2 currently achieves state-of-the-art results in unsupervised settings, we compare both paradigms to quantify the relative merits of label-free feature learning versus pseudo-label-driven supervision for scaling 3D segmentation models.
PointTransformerV3 (PTv3) [40]
@@
is a recently proposed method to accelerate transformer architectures and enable large-scale training by jointly training multiple datasets with diverse label spaces. This contrasts with the LabelMaker [37] approach, which translates label spaces to a common one before training. The combination of PTv3 and PPT achieves state-of-the-art performance on the ScanNet/ScanNet200 semantic segmentation benchmarks.
4.2 Datasets and Evaluation Metrics
Dataset #train #val #test real #classes